# Accelerating Ill-conditioned Hankel Matrix RecoveryHanQin Cai, Longxiu Huang, Xiliang Lu, and Juntao You

### v1

@@ 5.1 Technical lemmas
(31) $~{}\left\lvert\left\langle\Pi_{\widehat{\Omega}}\mathcal{G}^{*}\left({\bm{L}}_{k}{\bm{R}}_{k}^{*}-\mathcal{G}\bm{z}_{\star}\right),\mathcal{G}^{*}\left({\bm{L}}_{k}{\bm{R}}_{k}^{*}-\mathcal{G}\bm{z}_{\star}\right)\right\rangle\right\rvert^{\frac{1}{2}}$
(32) $\leq$ $~{}\left(1+\frac{c\kappa}{1-\varepsilon_{0}}\right)\sqrt{c_{s}\mu r|\widehat{\Omega}|/(2n)}\left(\left\lVert\bm{\Delta}_{{\bm{L}}_{k}}\bm{\Sigma}_{\star}^{1/2}\right\lVert_{\mathrm{F}}+\left\lVert\bm{\Delta}_{{\bm{R}}_{k}}\bm{\Sigma}_{\star}^{1/2}\right\lVert_{\mathrm{F}}\right),$
and for all ${\bm{L}}\in\mathbb{C}^{n_{1}\times r},~{}{\bm{R}}\in\mathbb{C}^{n_{2}\times r}$ it holds
@@
Proof 5.21.
By using the decomposition ${\bm{L}}_{k}{\bm{R}}_{k}^{*}-\mathcal{G}\bm{z}_{\star}={\bm{L}}_{k}{\bm{R}}_{k}^{*}-\mathcal{G}\bm{z}_{\star}=(\bm{L}_{\natural}-\bm{L}_{\star})\bm{R}_{\natural}^{*}+\bm{L}_{\star}(\bm{R}_{\natural}-\bm{R}_{\star})=(\bm{L}_{\natural}-\bm{L}_{\star})\bm{R}_{\star}^{*}+\bm{L}_{\natural}(\bm{R}_{\natural}-\bm{R}_{\star})$, we have
@@
where $\theta=\max\{\left\lVert\bm{R}_{\natural}\bm{\Sigma}_{\star}^{-1/2}\right\lVert_{2,\infty}+\left\lVert\bm{R}_{\star}\bm{\Sigma}_{\star}^{-1/2}\right\lVert_{2,\infty},\left\lVert\bm{L}_{\star}\bm{\Sigma}_{\star}^{-1/2}\right\lVert_{2,\infty}+\left\lVert\bm{L}_{\natural}\bm{\Sigma}_{\star}^{-1/2}\right\lVert_{2,\infty}\}\leq\left(1+\frac{c\kappa}{1-\varepsilon_{0}}\right)\sqrt{c_{s}\mu r/n}$ due to Lemma 5.16. Let $\widehat{\Omega}=\{\hat{a}_{k}\}_{1\leq k\leq|\widehat{\Omega}|}$, denote $\widehat{\Phi}_{k}:=\{\left(i,j\right):\left(i+j-1\right)=\hat{a}_{k}\}$ and $\widehat{\Phi}=\cup_{1\leq k\leq|\widehat{\Omega}|}\widehat{\Phi}_{k}$. Then, we have
$~{}\left\lvert\left\langle\Pi_{\widehat{\Omega}}\mathcal{G}^{*}\left({\bm{L}}_{k}{\bm{R}}_{k}^{*}-\mathcal{G}\bm{z}_{\star}\right),\mathcal{G}^{*}\left({\bm{L}}_{k}{\bm{R}}_{k}^{*}-\mathcal{G}\bm{z}_{\star}\right)\right\rangle\right\rvert=\sum_{\hat{a}_{k}\in\widehat{\Omega}}\frac{1}{\varsigma_{\hat{a}_{k}}}\left\lvert\sum_{\left(i,j\right)\in\widehat{\Phi}_{k}}[\bm{L}_{\natural}\bm{R}_{\natural}^{*}-\mathcal{G}\bm{z}_{\star}]_{i,j}\right\rvert^{2}$
$\leq$ $\sum_{\hat{a}_{k}\in\widehat{\Omega}}\sum_{\left(i,j\right)\in\widehat{\Phi}_{k}}\left\lvert[\bm{L}_{\natural}\bm{R}_{\natural}^{*}-\mathcal{G}\bm{z}_{\star}]^{2}_{i,j}\right\rvert=\sum_{\left(i,j\right)\in\widehat{\Phi}}\left\lvert[\bm{L}_{\natural}\bm{R}_{\natural}^{*}-\mathcal{G}\bm{z}_{\star}]^{2}_{i,j}\right\rvert$
@@
$\leq$ $~{}\frac{\theta^{2}}{2}\left(\sum_{j\in\widehat{\Phi}\left(:,i\right)}\sum_{i}\left\lVert[\bm{\Delta}_{{\bm{L}}_{k}}\bm{\Sigma}_{\star}^{1/2}]\left(i,:\right)\right\lVert_{2}^{2}+\sum_{i\in\widehat{\Phi}\left(:,j\right)}\sum_{j}\left\lVert[\bm{\Delta}_{{\bm{R}}_{k}}\bm{\Sigma}_{\star}^{1/2}]\left(j,:\right)\right\lVert_{2}^{2}\right)$
@@
where the first inequality follows from $\left\lvert\widehat{\Phi}_{k}\right\rvert=\varsigma_{\hat{a}_{k}}$ and the Cauchy-Schwarz inequality, the third inequality follows from $(a+b)^{2}\leq 2a^{2}+2b^{2}$, and the fourth inequality follow from $\mathrm{card}\left(\widehat{\Phi}\left(i,:\right)\right)\leq|\widehat{\Omega}|$, $\mathrm{card}\left(\widehat{\Phi}\left(:,j\right)\right)\leq|\widehat{\Omega}|$ due to the Hankel structure. Therefore, by $\sqrt{a^{2}+b^{2}}\leq\left\lvert a\right\rvert+\left\lvert b\right\rvert$ we conclude the inequality (31). Similarly, by
@@
we conclude the second statement.
Lemma 5.22.
@@
$~{}p^{-1}\left\lvert\left\langle\mathcal{G}\Pi_{\Omega}\left(\bm{s}-\bm{s}_{\star}\right),{\bm{L}}_{{\bm{A}}}{\bm{R}}_{{\bm{A}}}^{*}\right\rangle\right\rvert$
@@
where $\xi=\left((\gamma+1)\alpha\left(1+\frac{c\kappa}{1-\varepsilon_{0}}\right)\sqrt{\frac{c_{s}\mu r}{2}}+(2+2\varepsilon_{0})\frac{\sqrt{(\gamma+1)\alpha}}{\sqrt{(\gamma-1)}}\right).$
Proof 5.23.
@@
(33) $\left\lvert\left[\Pi_{\Omega}\bm{s}_{\star}-\bm{s}\right]_{i}\right\rvert=\left\lvert\left[\Pi_{\Omega}\left(\bm{z}_{\star}-\mathcal{G}^{*}\left({\bm{L}}_{k}{\bm{R}}_{k}^{*}\right)\right)\right]_{i}\right\rvert=\left\lvert\left[\Pi_{\Omega}(\mathcal{G}^{*}\bm{\Delta}_{k})\right]_{i}\right\rvert,~{}\text{for}~{}i\in\Omega^{s},$
@@
(34) $\left\lvert\left[\Pi_{\Omega}\bm{s}_{\star}\right]_{i}\right\rvert\leq{\frac{\left\lvert\left\langle\Pi_{\Omega}\left(\mathcal{G}^{*}\bm{\Delta}_{k}\right),\mathcal{G}^{*}\bm{\Delta}_{k}\right\rangle\right\rvert^{1/2}}{\sqrt{(\gamma-1)\alpha m}}}+\left\lvert\left[\Pi_{\Omega}\left(\mathcal{G}^{*}\bm{\Delta}_{k}\right)\right]_{i}\right\rvert,~{}\text{for}~{}i\in\Omega^{s_{\star}}\setminus\Omega^{s}.$
@@
$\leq$ $~{}\sum_{i\in\Omega^{s}\cup\Omega^{s_{\star}}}\left\lvert\left[\Pi_{\Omega}\left(\mathcal{G}^{*}\bm{\Delta}_{k}\right)\right]_{i}\right\rvert\left\lvert\left[\mathcal{G}^{*}({\bm{L}}_{{\bm{A}}}{\bm{R}}_{{\bm{A}}}^{*})\right]_{i}\right\rvert$
$~{}+\sum_{i\in\Omega^{s_{\star}}\setminus\Omega^{s}}{\frac{\left\lvert\left\langle\Pi_{\Omega}\left(\mathcal{G}^{*}\bm{\Delta}_{k}\right),\mathcal{G}^{*}\bm{\Delta}_{k}\right\rangle\right\rvert^{\frac{1}{2}}}{\sqrt{(\gamma-1)\alpha m}}\left\lvert\left[\mathcal{G}^{*}({\bm{L}}_{{\bm{A}}}{\bm{R}}_{{\bm{A}}}^{*})\right]_{i}\right\rvert}$
@@
where the third inequality follows form $ab\leq\frac{\beta}{2}a^{2}+\frac{b^{2}}{2\beta}$ for any $\beta>0$, and the Cauchy-Schwarz inequality, the fourth inequality follows from Lemma 5.20. Letting
@@
where $\xi=\left((\gamma+1)\alpha\left(1+\frac{c\kappa}{1-\varepsilon_{0}}\right)\sqrt{\frac{c_{s}\mu r}{2}}+(2+2\varepsilon_{0})\frac{\sqrt{(\gamma+1)\alpha}}{\sqrt{(\gamma-1)}}\right)$ and the last inequality follows from Lemma 5.18.
@@ 5.4.1 Contraction Property in Full Observation
$\left\lVert\left(\widetilde{{\bm{L}}}_{k+1}{\bm{Q}}_{k}-\bm{L}_{\star}\right)\bm{\Sigma}_{\star}^{1/2}\right\lVert_{\mathrm{F}}^{2}=\left\lVert\bm{T}_{1}\right\lVert_{\mathrm{F}}^{2}+\left\lVert\bm{T}_{2}\right\lVert_{\mathrm{F}}^{2}-2\mathrm{Re}\left\langle\bm{T}_{1},\bm{T}_{2}\right\rangle,$
@@
Therefore, if provided $\alpha\leq\frac{\varepsilon^{2}_{0}(\gamma-1)}{64(\gamma+1)(1+c\kappa)^{2}c_{s}\mu r}$, combing all the pieces above we have
@@
Then, Lemma 5.12 gives $d_{k+1}\leq(1-0.6\eta)d_{k}$ and $({\bm{L}}_{k+1},{\bm{R}}_{k+1})\in\mathcal{E}(\varepsilon_{0}\sigma^{\star}_{r},c\sqrt{\frac{c_{s}\mu r}{n}}\sigma^{\star}_{1})$. Finally, we complete the proof by proving Claim 1.
Proof 5.25 (Proof of Claim 1).
The upper bound of $\left\lVert\bm{T}_{1}\right\lVert_{\mathrm{F}}$ following the argument as in [33, Eq. (46)]. For $\left\lvert\left\langle\bm{T}_{1},\bm{T}_{2}\right\rangle\right\rvert$, we denote
(35) $\begin{split}{\bm{L}}_{{\bm{A}}_{1}}&:=\bm{\Delta}_{{\bm{L}}_{k}}\bm{\Sigma}_{\star}\left(\bm{R}_{\natural}^{*}\bm{R}_{\natural}\right)^{-1}\bm{\Sigma}_{\star}^{1/2},~{}{\bm{R}}_{{\bm{A}}_{1}}:=\bm{R}_{\natural}\bm{\Sigma}_{\star}^{-1/2},\cr{\bm{L}}_{{\bm{A}}_{2}}&:=\bm{L}_{\star}\bm{\Sigma}_{\star}^{-1/2}={\bm{U}}_{\star},~{}{\bm{R}}_{{\bm{A}}_{2}}:=\bm{R}_{\natural}\left(\bm{R}_{\natural}^{*}\bm{R}_{\natural}\right)^{-1}\bm{\Sigma}_{\star}\left(\bm{R}_{\natural}^{*}\bm{R}_{\natural}\right)^{-1}\bm{R}_{\natural}^{*}\bm{\Delta}_{{\bm{R}}_{k}}\bm{\Sigma}_{\star}^{1/2}.\end{split}$
Then, by applying Lemma 5.22 with $m=n$ we have
@@
(36) $\left\lVert\bm{\Sigma}_{\star}^{1/2}\left(\bm{R}_{\natural}^{*}\bm{R}_{\natural}\right)^{-1}\bm{\Sigma}_{\star}^{1/2}\right\lVert_{2}$ $=\left\lVert\bm{R}_{\natural}\left(\bm{R}_{\natural}^{*}\bm{R}_{\natural}\right)^{-1}\bm{\Sigma}_{\star}^{1/2}\right\lVert_{2}^{2},$
(37) $\left\lVert\bm{R}_{\natural}\left(\bm{R}_{\natural}^{*}\bm{R}_{\natural}\right)^{-1}\bm{\Sigma}_{\star}\left(\bm{R}_{\natural}^{*}\bm{R}_{\natural}\right)^{-1}\bm{R}_{\natural}^{*}\right\lVert_{2}$ $=\left\lVert\bm{R}_{\natural}\left(\bm{R}_{\natural}^{*}\bm{R}_{\natural}\right)^{-1}\bm{\Sigma}_{\star}^{1/2}\right\lVert_{2}^{2},$
(38) (29d)$\left\lVert{\bm{L}}_{{\bm{A}}_{1}}\right\lVert_{\mathrm{F}}\leq\frac{1}{(1-\varepsilon_{0})^{2}}\left\lVert\bm{\Delta}_{{\bm{L}}_{k}}\bm{\Sigma}_{\star}^{1/2}\right\lVert_{\mathrm{F}},$ $~{}\left\lVert{\bm{R}}_{{\bm{A}}_{2}}\right\lVert_{\mathrm{F}}\leq\frac{1}{(1-\varepsilon_{0})^{2}}\left\lVert\bm{\Delta}_{{\bm{L}}_{k}}\bm{\Sigma}_{\star}^{1/2}\right\lVert_{\mathrm{F}},$
@@
where the first inequality follows from Lemma 5.22, the last inequality follows from (29c) and similar argument to (36).

### v2

@@ 5.4 Proof of Theorems 2 and 3 (local convergence)
$~{}\left\lvert\left\langle\Pi_{\widehat{\Omega}}\mathcal{G}^{*}\left({\bm{L}}_{k}{\bm{R}}_{k}^{*}-\mathcal{G}\bm{z}_{\star}\right),\mathcal{G}^{*}\left({\bm{L}}_{k}{\bm{R}}_{k}^{*}-\mathcal{G}\bm{z}_{\star}\right)\right\rangle\right\rvert^{\frac{1}{2}}$ (30)
$\leq$ $~{}\left(1+\frac{c\kappa}{1-\varepsilon_{0}}\right)\sqrt{c_{s}\mu r|\widehat{\Omega}|/(2n)}\left(\left\lVert\bm{\Delta}_{{\bm{L}}_{k}}\bm{\Sigma}_{\star}^{1/2}\right\lVert_{\mathrm{F}}+\left\lVert\bm{\Delta}_{{\bm{R}}_{k}}\bm{\Sigma}_{\star}^{1/2}\right\lVert_{\mathrm{F}}\right),$ (31)
and for all ${\bm{L}}\in\mathbb{C}^{n_{1}\times r},~{}{\bm{R}}\in\mathbb{C}^{n_{2}\times r}$ it holds that
@@
Proof.
By using the decomposition ${\bm{L}}_{k}{\bm{R}}_{k}^{*}-\mathcal{G}\bm{z}_{\star}=(\bm{L}_{\natural}-\bm{L}_{\star})\bm{R}_{\natural}^{*}+\bm{L}_{\star}(\bm{R}_{\natural}-\bm{R}_{\star})=(\bm{L}_{\natural}-\bm{L}_{\star})\bm{R}_{\star}^{*}+\bm{L}_{\natural}(\bm{R}_{\natural}-\bm{R}_{\star})$, we have
@@
where
$\theta$ $=\max\{\left\lVert\bm{R}_{\natural}\bm{\Sigma}_{\star}^{-1/2}\right\lVert_{2,\infty}+\left\lVert\bm{R}_{\star}\bm{\Sigma}_{\star}^{-1/2}\right\lVert_{2,\infty},\left\lVert\bm{L}_{\star}\bm{\Sigma}_{\star}^{-1/2}\right\lVert_{2,\infty}+\left\lVert\bm{L}_{\natural}\bm{\Sigma}_{\star}^{-1/2}\right\lVert_{2,\infty}\}$
$\leq\left(1+\frac{c\kappa}{1-\varepsilon_{0}}\right)\sqrt{c_{s}\mu r/n}$
due to Lemma 8. Let $\widehat{\Omega}=\{\hat{a}_{t}\}_{1\leq t\leq|\widehat{\Omega}|}$, denote $\widehat{\Phi}_{t}:=\{\left(i,j\right):\left(i+j-1\right)=\hat{a}_{t}\}$ and $\widehat{\Phi}=\cup_{1\leq t\leq|\widehat{\Omega}|}\widehat{\Phi}_{t}$. Then, by definition in (9), we have
$~{}\left\lvert\left\langle\Pi_{\widehat{\Omega}}\mathcal{G}^{*}\left({\bm{L}}_{k}{\bm{R}}_{k}^{*}-\mathcal{G}\bm{z}_{\star}\right),\mathcal{G}^{*}\left({\bm{L}}_{k}{\bm{R}}_{k}^{*}-\mathcal{G}\bm{z}_{\star}\right)\right\rangle\right\rvert=\sum_{\hat{a}_{t}\in\widehat{\Omega}}\frac{1}{\varsigma_{\hat{a}_{t}}}\left\lvert\sum_{\left(i,j\right)\in\widehat{\Phi}_{t}}[\bm{L}_{\natural}\bm{R}_{\natural}^{*}-\mathcal{G}\bm{z}_{\star}]_{i,j}\right\rvert^{2}$
$\leq$ $\sum_{\hat{a}_{t}\in\widehat{\Omega}}\sum_{\left(i,j\right)\in\widehat{\Phi}_{t}}\left\lvert[\bm{L}_{\natural}\bm{R}_{\natural}^{*}-\mathcal{G}\bm{z}_{\star}]^{2}_{i,j}\right\rvert=\sum_{\left(i,j\right)\in\widehat{\Phi}}\left\lvert[\bm{L}_{\natural}\bm{R}_{\natural}^{*}-\mathcal{G}\bm{z}_{\star}]^{2}_{i,j}\right\rvert$
@@
$\leq$ $~{}\sum_{\left(i,j\right)\in\widehat{\Phi}}\frac{\theta^{2}}{2}\left(\left\lVert[\bm{\Delta}_{{\bm{L}}_{k}}\bm{\Sigma}_{\star}^{1/2}]\left(i,:\right)\right\lVert_{2}^{2}+\left\lVert[\bm{\Delta}_{{\bm{R}}_{k}}\bm{\Sigma}_{\star}^{1/2}]\left(j,:\right)\right\lVert_{2}^{2}\right)$
$=$ $~{}\frac{\theta^{2}}{2}\Bigg{(}\sum_{i}\sum_{1\leq t\leq|\hat{\Omega}|}\chi_{[1,\infty)}(\hat{a}_{t}+1-i)\left\lVert[\bm{\Delta}_{{\bm{L}}_{k}}\bm{\Sigma}_{\star}^{1/2}]\left(i,:\right)\right\lVert_{2}^{2}+\sum_{j}\sum_{1\leq t\leq|\hat{\Omega}|}\chi_{[1,\infty)}(\hat{a}_{t}+1-j)\left\lVert[\bm{\Delta}_{{\bm{R}}_{k}}\bm{\Sigma}_{\star}^{1/2}]\left(j,:\right)\right\lVert_{2}^{2}\Bigg{)}$
$\leq$ $~{}\frac{\theta^{2}}{2}\left(\sum_{i}|\hat{\Omega}|\left\lVert[\bm{\Delta}_{{\bm{L}}_{k}}\bm{\Sigma}_{\star}^{1/2}]\left(i,:\right)\right\lVert_{2}^{2}+\sum_{j}|\hat{\Omega}|\left\lVert[\bm{\Delta}_{{\bm{R}}_{k}}\bm{\Sigma}_{\star}^{1/2}]\left(j,:\right)\right\lVert_{2}^{2}\right)$
@@
where the first inequality follows from $\left\lvert\widehat{\Phi}_{t}\right\rvert=\varsigma_{\hat{a}_{t}}$ and the Cauchy-Schwarz inequality, the third inequality follows from $(a+b)^{2}\leq 2a^{2}+2b^{2}$, and $\chi_{[1,\infty)}$ is the indicator function of $[1,\infty)$. Therefore, by $\sqrt{a^{2}+b^{2}}\leq\left\lvert a\right\rvert+\left\lvert b\right\rvert$ we conclude the inequality (30). Similarly, by
@@
we conclude the second statement. ∎
Lemma 11.
@@
$~{}\frac{1}{p}\left\lvert\left\langle\mathcal{G}\Pi_{\Omega}\left(\bm{s}-\bm{s}_{\star}\right),{\bm{L}}_{{\bm{A}}}{\bm{R}}_{{\bm{A}}}^{*}\right\rangle\right\rvert$
@@
where $\xi=\left((\gamma+1)\alpha\left(1+\frac{c\kappa}{1-\varepsilon_{0}}\right)\sqrt{\frac{c_{s}\mu r}{2}}+(1+3\sqrt{\varepsilon_{0}})\frac{\sqrt{(\gamma+1)\alpha}}{\sqrt{(\gamma-1)}}\right)$.
Proof.
@@
$\left\lvert\left[\Pi_{\Omega}\bm{s}_{\star}-\bm{s}\right]_{i}\right\rvert=\left\lvert\left[\Pi_{\Omega}\left(\bm{z}_{\star}-\mathcal{G}^{*}\left({\bm{L}}_{k}{\bm{R}}_{k}^{*}\right)\right)\right]_{i}\right\rvert=\left\lvert\left[\Pi_{\Omega}(\mathcal{G}^{*}\bm{\Delta}_{k})\right]_{i}\right\rvert,~{}\text{for}~{}i\in\Omega^{s},$ (32)
@@
$\left\lvert\left[\Pi_{\Omega}\bm{s}_{\star}\right]_{i}\right\rvert\leq{\frac{\left\lvert\left\langle\Pi_{\Omega}\left(\mathcal{G}^{*}\bm{\Delta}_{k}\right),\mathcal{G}^{*}\bm{\Delta}_{k}\right\rangle\right\rvert^{1/2}}{\sqrt{(\gamma-1)\alpha m}}}+\left\lvert\left[\Pi_{\Omega}\left(\mathcal{G}^{*}\bm{\Delta}_{k}\right)\right]_{i}\right\rvert,~{}\text{for}~{}i\in\Omega^{s_{\star}}\setminus\Omega^{s}.$ (33)
@@
$\leq$ $~{}\sum_{i\in\Omega^{s}\cup\Omega^{s_{\star}}}\left\lvert\left[\Pi_{\Omega}\left(\mathcal{G}^{*}\bm{\Delta}_{k}\right)\right]_{i}\right\rvert\left\lvert\left[\mathcal{G}^{*}({\bm{L}}_{{\bm{A}}}{\bm{R}}_{{\bm{A}}}^{*})\right]_{i}\right\rvert+\sum_{i\in\Omega^{s_{\star}}\setminus\Omega^{s}}{\frac{\left\lvert\left\langle\Pi_{\Omega}\left(\mathcal{G}^{*}\bm{\Delta}_{k}\right),\mathcal{G}^{*}\bm{\Delta}_{k}\right\rangle\right\rvert^{\frac{1}{2}}}{\sqrt{(\gamma-1)\alpha m}}\left\lvert\left[\mathcal{G}^{*}({\bm{L}}_{{\bm{A}}}{\bm{R}}_{{\bm{A}}}^{*})\right]_{i}\right\rvert}$
@@
where the third inequality follows form $ab\leq\frac{\beta}{2}a^{2}+\frac{b^{2}}{2\beta}$ for any $\beta>0$, and the Cauchy-Schwarz inequality, the fourth inequality follows from Lemma 10. Letting
@@
where $\xi=\left((\gamma+1)\alpha\left(1+\frac{c\kappa}{1-\varepsilon_{0}}\right)\sqrt{\frac{c_{s}\mu r}{2}}+(1+3\sqrt{\varepsilon_{0}})\frac{\sqrt{(\gamma+1)\alpha}}{\sqrt{(\gamma-1)}}\right)$ and the last inequality follows from Lemma 9. ∎
@@ 5.4.1 Contraction Property in Full Observation
$\left\lVert\left(\widetilde{{\bm{L}}}_{k+1}{\bm{Q}}_{k}-\bm{L}_{\star}\right)\bm{\Sigma}_{\star}^{1/2}\right\lVert_{\mathrm{F}}^{2}=\left\lVert\bm{T}_{1}\right\lVert_{\mathrm{F}}^{2}+\left\lVert\bm{T}_{2}\right\lVert_{\mathrm{F}}^{2}-2\mathrm{Re}\left\langle\bm{T}_{1},\bm{T}_{2}\right\rangle,$ (34)
@@
Recall that $\xi=\left((\gamma+1)\alpha\left(1+\frac{c\kappa}{1-\varepsilon_{0}}\right)\sqrt{\frac{c_{s}\mu r}{2}}+(1+3\sqrt{\varepsilon_{0}})\frac{\sqrt{(\gamma+1)\alpha}}{\sqrt{(\gamma-1)}}\right)$. Therefore, provided $\alpha\leq\frac{\varepsilon^{2}_{0}(\gamma-1)}{64(\gamma+1)(1+c\kappa)^{2}c_{s}\mu r}$, by substituting the above three inequalities in Claim 1 to (34), we have
@@
Then, Lemma 7 gives $d_{k+1}\leq(1-0.6\eta)d_{k}$ and $({\bm{L}}_{k+1},{\bm{R}}_{k+1})\in\mathcal{E}(\varepsilon_{0}\sigma^{\star}_{r},c\sqrt{\frac{c_{s}\mu r}{n}}\sigma^{\star}_{1})$. Finally, we will complete the proof by proving Claim 1. ∎
Proof of Claim 1.
The upper bound of $\left\lVert\bm{T}_{1}\right\lVert_{\mathrm{F}}$ follows the argument as in [36, Eq. (46)]. For $\left\lvert\left\langle\bm{T}_{1},\bm{T}_{2}\right\rangle\right\rvert$, we denote
$\begin{split}{\bm{L}}_{{\bm{A}}_{1}}&:=\bm{\Delta}_{{\bm{L}}_{k}}\bm{\Sigma}_{\star}\left(\bm{R}_{\natural}^{*}\bm{R}_{\natural}\right)^{-1}\bm{\Sigma}_{\star}^{1/2},~{}{\bm{R}}_{{\bm{A}}_{1}}:=\bm{R}_{\natural}\bm{\Sigma}_{\star}^{-1/2},\cr{\bm{L}}_{{\bm{A}}_{2}}&:=\bm{L}_{\star}\bm{\Sigma}_{\star}^{-1/2}={\bm{U}}_{\star},~{}{\bm{R}}_{{\bm{A}}_{2}}:=\bm{R}_{\natural}\left(\bm{R}_{\natural}^{*}\bm{R}_{\natural}\right)^{-1}\bm{\Sigma}_{\star}\left(\bm{R}_{\natural}^{*}\bm{R}_{\natural}\right)^{-1}\bm{R}_{\natural}^{*}\bm{\Delta}_{{\bm{R}}_{k}}\bm{\Sigma}_{\star}^{1/2}.\end{split}$ (35)
Then, by applying Lemma 11 with $m=n$ we have
@@
$\left\lVert\bm{\Sigma}_{\star}^{1/2}\left(\bm{R}_{\natural}^{*}\bm{R}_{\natural}\right)^{-1}\bm{\Sigma}_{\star}^{1/2}\right\lVert_{2}$ $=\left\lVert\bm{R}_{\natural}\left(\bm{R}_{\natural}^{*}\bm{R}_{\natural}\right)^{-1}\bm{\Sigma}_{\star}^{1/2}\right\lVert_{2}^{2},$ (36)
$\left\lVert\bm{R}_{\natural}\left(\bm{R}_{\natural}^{*}\bm{R}_{\natural}\right)^{-1}\bm{\Sigma}_{\star}\left(\bm{R}_{\natural}^{*}\bm{R}_{\natural}\right)^{-1}\bm{R}_{\natural}^{*}\right\lVert_{2}$ $=\left\lVert\bm{R}_{\natural}\left(\bm{R}_{\natural}^{*}\bm{R}_{\natural}\right)^{-1}\bm{\Sigma}_{\star}^{1/2}\right\lVert_{2}^{2},$ (37)
(28d)$\left\lVert{\bm{L}}_{{\bm{A}}_{1}}\right\lVert_{\mathrm{F}}\leq\frac{1}{(1-\varepsilon_{0})^{2}}\left\lVert\bm{\Delta}_{{\bm{L}}_{k}}\bm{\Sigma}_{\star}^{1/2}\right\lVert_{\mathrm{F}},$ $~{}\left\lVert{\bm{R}}_{{\bm{A}}_{2}}\right\lVert_{\mathrm{F}}\leq\frac{1}{(1-\varepsilon_{0})^{2}}\left\lVert\bm{\Delta}_{{\bm{L}}_{k}}\bm{\Sigma}_{\star}^{1/2}\right\lVert_{\mathrm{F}},$ (38)
@@
where the first inequality follows from Lemma 11, the last inequality follows from (28c) and similar argument to (36). ∎